# On the parameterized complexity of computing tree-partitions

### v4

@@ Abstract
Parameterized algorithms; Tree-partitions; Tree-partition-width; Treewidth; Domino Treewidth; Approximation Algorithms; Parameterized Complexity
@@ 1 Introduction
Graph decompositions have been a very useful tool to draw the line between tractability and intractability of computational problems. There are many meta-theorems showing that a collection of problems can be solved efficiently if a decomposition of some form is given (e.g. for treewidth [Cou90], for clique-width [CMR00], for twin-width [BKTW22], for mim-width [BDJ23]). By finding efficient algorithms to compute a decomposition if it exists, we deduce the existence of efficient algorithms even if the decomposition is not given. In particular, this proves useful when designing win-win arguments: for some problems, the existence of a solution and the existence of a decomposition are not independent, so that we can either use the decomposition for an efficient computation of the solution, or conclude that a solution must (or cannot) exist when there is no decomposition of small enough width.
The most successful notion of graph decomposition to date is certainly tree decompositions, and its corresponding parameter treewidth. Any problem expressible in MSO2111Formulae with quantification over sets of edges or vertices, quantification over vertices and edges, and with the incidence predicate. can be solved in linear time in graphs of bounded treewidth due to a meta-theorem of Courcelle [Cou90] and the algorithm of Bodlaender for computing an optimal tree decomposition [Bod96]. Treewidth is a central tool in the study of minor-closed graph classes. A minor-closed graph class has bounded treewidth if and only if it contains no large grid minor.
In this paper, we focus on the parameter tree-partition-width (also called strong treewidth) which was independently introduced by Seese [See85] and Halin [Hal91]. It is known to have simple relations to treewidth [DO95, Woo09]: $\operatorname{\textup{{tw}}}=O(\operatorname{\textup{{tpw}}})$, and $\operatorname{\textup{{tpw}}}=O(\Delta\operatorname{\textup{{tw}}})$, where $\operatorname{\textup{{tw}}},\operatorname{\textup{{tpw}}},\Delta$ denote the treewidth, the tree-partition-width, and the maximum degree respectively. Applications of tree-partition-width include graph drawing and graph colouring [CDMW08, GLM05, DMW05, DSW07, WT06, BW08, ADOV03]. Recently, Bodlaender, Cornelissen and Van der Wegen [BCvdW22] showed for a number of problems (in particular, problems related to network flow) that these are intractable (XNLP-complete) when the pathwidth is used as parameter, but become fixed parameter tractable when parameterized by the width of a given tree-partition. This raises the question of the complexity of finding tree-partitions. We show that computing tree-partitions of approximate width is tractable.
@@
Thus, this removes the requirement from the results from [BCvdW22] that a tree-partition of small width is part of the input. Our technique is modular and allows us to also give alternatives running in FPT time or polynomial time with an improved approximation factor (see Theorem 3.11). Although not formulated as an algorithm, a construction of Ding and Oporowski [DO96] implies an FPT algorithm to compute tree-partitions of width $f(k)$ for graphs of tree-partition-width $k$, for some fixed computable function $f$. We adapt their construction and give some new arguments designed for our purposes. This significantly improves on the upper bounds to the width, and the running time.
The results from [BCvdW22] are stated in terms of the notions of stable gonality, stable tree-partition-width and a new parameter called weighted tree-partition-width [^2] In earlier versions of [BCvdW22], the parameter was called treebreadth, but to avoid confusion, the term weighted tree-partition-width is now used.. The notion of stability comes from the algebraic geometry origins of the notion of gonality; in graph-theoretic terms, this implies that we look at the minimum over all possible subdivisions of edges. It turns out that tree-partition-width, stable tree-partition-width, and weighted tree-partition-width (with edge weights one) are bounded by polynomial functions of each other; see Section 5 and Corollary 5.7. In Section 6, we obtain some results on the complexity of computing and approximating the weighted tree-partition-width, as corollaries of earlier results.
Related to tree-partition-width is the notion of domino treewidth, first studied by Bodlaender and Engelfriet [BE97]. A domino tree decomposition is a tree decomposition where each vertex is in at most two bags. Where graphs of small tree-partition-width can have large degree, a graph of domino treewidth $k$ has maximum degree at most $2k$. Bodlaender and Engelfriet show that Domino Treewidth is hard for each class $W[t]$, $t\in\mathbb{N}$; we improve this result and show XALP-completeness.
@@
We also consider the parameter tree-cut width introduced by Wollan in [Wol15]. As the tractability results of Bodlaender et al. [BCvdW22] use techniques similar to a previous work on algorithmic applications of tree-cut width [GKS15], one may wonder whether there is a relationship between tree-cut width and tree-partition-width.
@@
  • We obtain a parameter that is polynomially tied to tree-partition-width and is topological minor monotone (see Theorem 5.6). We use this to show that tree-partition-width is relatively stable with respect to subdivisions: if we define $\underline{\operatorname{\textup{{tpw}}}}(G)$ (resp. $\overline{\operatorname{\textup{{tpw}}}}(G)$) as the minimum (resp. maximum) tree-partition-width over subdivisions of $G$, then $,\leavevmode\nobreak\ \operatorname{\textup{{tpw}}}$ tpw and $\overline{\operatorname{\textup{{tpw}}}}$ are polynomially tied. The parameter $\underline{\operatorname{\textup{{tpw}}}}(G)$ corresponds to ‘stable tree-partition-width’.
@@
  • On the other hand, a bound on tree-partition-width does not imply a bound on tree-cut width (see Observation 5).
@@ 2 Preliminaries
We also consider a variant of the notion for weighted graphs. The notion was introduced in [BCvdW22]. We use a slight generalization where we allow also weights of vertices, to facilitate some of our proofs.
@@ 3.1 Description of the algorithm
As mentioned above, we do not directly specify the function $w=w(k)$, since different algorithms for step 1 give different solution qualities (bounds for $w(k)$) and running times. Since $\operatorname{\textup{{tw}}}+1\leq 2\cdot\operatorname{\textup{{tpw}}}$, if $\operatorname{\textup{{tw}}}(G)>2k-1$ it follows that $\operatorname{\textup{{tpw}}}(G)>k$. If we obtained a decomposition of width $w$, we also know that $\operatorname{\textup{{tw}}}(G)\leq 2k-1$, and hence there are at most $(2k-1)n$ edges in $G$.
@@
We later describe several ways of computing the edges of $G^{b}$. We will show in Claim 1 that vertices in the same connected component of $G^{b}$ must be in the same bag for any tree-partition of width at most $k$. For this reason, we conclude that $\operatorname{\textup{{tpw}}}(G)>k$ if a component of $G^{b}$ has more than $k$ vertices.
@@
Given the components of $G^{b}$, we can compute $H$, and its 2-connected components in time $O(k^{O(1)}n)$. Using Claim 2, we obtain a tree decomposition of $H$ by replacing vertices of $G$, in the tree decomposition of $G$, by their component in $G^{b}$.
By Claim 3, the maximum degree $\Delta_{H}$ within the 2-connected components of $H$ is at most $Cbk^{2}$ for some constant $C$ when $\operatorname{\textup{{tpw}}}(G)\leq k$.
@@
By rooting the decomposition of $H$ in 2-connected components, we can define a parent cutvertex for each 2-connected component except the root. We separately compute tree partitions for each 2-connected component of $H$ with the constraint that their parent cutvertex should be the single vertex of its bag. A construction of Wood [Woo09] enables us to compute a tree-partition of width $O(\Delta w)$ for any graph of maximum degree $\Delta$ and treewidth $w$; this can be adjusted to allow for this isolation constraint without increasing the upper bound on the width. We give the details of this in Corollary 3.5. After doing this, the partitions of each component can be combined without increasing the width. Indeed, although cutvertices are shared, only one 2-connected component will consider putting other vertices in its bag. We obtain a tree-partition of $H$ of width $O(wbk^{2})$.
@@ 3.2 Correctness
Assume that $s$ and $t$ are in adjacent bags but not in the same bag for some tree-partition of width at most $k$. We denote their respective bags by $B_{s}$ and $B_{t}$. Then, $(B_{s}\cup B_{t})\setminus\{s,t\}$ is an $s$-$t$ separator of $G-st$. Consequently, $\mu(s,t)\leq 2k-2$. This proves the second point.
Claim 2.
@@
Proof 3.2.
Every component of $G^{b}$ appears in at least one $Y_{i}$, because it contains a vertex which must appear in at least one $X_{i}$. Furthermore, for each edge $UV$ of $H$, there must be vertices $u\in U,v\in V$ such that $uv$ is an edge of $G$. Hence, there is a bag $X_{i}$ containing $u$ and $v$ so $Y_{i}$ contains $U$ and $V$. Finally, suppose that there is a bag $Y_{i}$ not containing a component $C$ of $G^{b}$, and several components of $T-i$ have bags containing $C$. There must be an edge of $G^{b}$ connecting vertices $u$ and $v$ of $C$ such that $u$ is in bags of $X$ and $v$ is in bags of $X^{\prime}$, where $X$ and $X^{\prime}$ are in different components of $T-i$. By definition of $G^{b}$, the minimal size of a separator of $u$ and $v$ in $G$ is at least $b\geq w+1$. However, since the tree decomposition $(T,(X_{i}))$ has width $w$ and the bags containing $u$ are disjoint from the bags containing $v$ (in particular $X_{i}$ separates them), there is a separator of $u$ and $v$ of size at most $w$, a contradiction. This concludes the proof that $(T,(Y_{i}))$ is a tree decomposition of $H$.
Claim 3.
@@
Proof 3.3.
Consider $u$ a vertex achieving maximum degree in $B$. By definition of $B$, $B-u$ is connected. We denote by $N$ the neighbourhood of $u$ in $B$. Let $T$ be a spanning tree of $B-u$. We iteratively remove leaves that are not in $N$. This produces the reduced tree $T^{\prime}$. The maximum degree in this tree is $b-1$ as the set of edges incident to a given vertex can be extended to disjoint paths leading to vertices in $N$, hence leading to $u$. Since $H$ is a $b$-reduction, the number of disjoint paths between two vertices must be less than $b$.
Clearly the neighbours of $u$ must be either in the same bag as $u$ or in a neighbouring bag. Since the bag of $u$ will be a separator of vertices that are in distinct neighbouring bags, in particular, it splits the graph into several components each containing at most $k$ neighbours of $u$.
There must exist a subset of vertices of $T^{\prime}$ of size at most $k-1$ whose removal splits $T^{\prime}$ in components containing vertices of $N$ of total weight at most $k$. Since the degree of a vertex of $T^{\prime}$ is at most $b-1$, removing one of its vertices adds at most $b-2$ new components. Hence, after removing $k-1$ vertices, there are at most $1+(k-1)(b-2)$ components. We conclude that $|N|\leq k(1+(k-1)(b-2))$. Since $u$ had maximum degree in $B$, we conclude.
In [Woo09], Wood shows the following lemma.
@@
We deduce this slightly stronger version of [Woo09, Theorem 1]
@@
Proof 3.6.
We wish to apply Lemma 3.4 to $S\supseteq N(v)$. Let $\gamma=1+\sqrt{2}$. We have $|N(v)|\leq\Delta$, so in particular, $|N(v)|\leq 3(\gamma+1)(w+1)\Delta$. In case $|N(v)|<(\gamma+1)(w+1)$, we can add arbitrary vertices to $N(v)$ to form $S$ satisfying $|S|\geq(\gamma+1)(w+1)$. Otherwise, we simply set $S=N(v)$. We then apply the lemma to $S$ in $G-v$. There is a single bag that contains $N(v)$, and so we may add the bag $\{v\}$ adjacent to this in order to deduce a tree partition of $G$ of width $O(\Delta w)$ in which $v$ is the only vertex of its bag.
@@ 3.3 Time/quality trade-offs
  • An algorithm of Korhonen [Kor21] computes a tree decomposition of width at most $12k+$ 1 or reports that $\operatorname{\textup{{tw}}}(G)>k$ in time $2^{O(k)}n$.
  • An algorithm of Fomin et al. [FLS+18] computes a tree decomposition of width $O(k^{2})$ or reports that $\operatorname{\textup{{tw}}}(G)>k$ in time $O(k^{7}n\log n)$.
  • An algorithm of Feige et al. [FHL08] computes a tree decomposition of width $O(k\sqrt{\log k})$ or reports that $\operatorname{\textup{{tw}}}(G)>k$ in time $O(n^{O(1)})$.
@@
  • We can use a maximum-flow algorithm (e.g. Ford-Fulkerson [FF56]) to compute for each pair $\{s,t\}$ of vertices of $G$ whether there are at least $b$ vertex disjoint paths from $s$ to $t$, in time $O(bkn)$. To compute a minimum vertex cut, replace each vertex $v$ by two vertices $v_{\text{in}},v_{\text{out}}$ with an arc from $v_{\text{in}}$ to $v_{\text{out}}$. All arcs going to $v$ should go to $v_{\text{in}}$, and all arcs leaving $v$ should leave $v_{\text{out}}$. All arcs are given capacity 1. We may stop the maximum flow algorithm as soon as a flow of at least $b$ was found. Furthermore, we can reduce the number of pairs $\{s,t\}$ of vertices to check to $O(wn)$, as each pair must be contained in a bag due to $b\geq w+1$. This results in a total time of $O(wbkn^{2})$.
  • We can also use dynamic programming to enumerate all possible ways of connecting pairs of vertices that are in the same bag in time $2^{O(w\log w)}n$, which is sufficient to compute $G^{b}$. A state of the dynamic programming consists of the subset of vertices of the bag that are used by the partial solution, a matching on some of these vertices, up to two vertices that were decided as endpoints of the constructed paths, the number of already constructed paths between the endpoints, and two disjoint subsets of the used vertices that are not endpoints, nor in the matching such that we found a disjoint path from the first or second endpoint to them. The bound of $2^{O(w\log w)}$ follows from the fact that this is a subset of the labelled forests on $w$ vertices. We may assume that our tree decomposition is rooted and binary. We first tabulate answers for each subtree of the decomposition by starting from the leaves, and then tabulate answers for each complement of a subtree by starting from the root and, when branching to some child, combining with the partial solutions of the subtree of the other child. By combining tabulated values for subtrees and their complements, we obtain the sought information.
The $b$-reduction $H$ of $G$ and its 2-connected components can be computed in $O(k^{O(1)}n)$ time (see e.g. [HT73]), since the size of the graph is $O(k^{O(1)}n)$ here.
We will now make use of the following result due to Bodlaender and Hagerup [BH98]:
@@
When implementing the construction of Wood for 2-connected components of $H$, the running time is dominated by $O(n)$ queries to find a balanced separator with respect to a set $S$ of size $k^{O(1)}$. After a preprocessing in time $O(kn)$, we can do this in time $k^{O(1)}d$ where $d$ is the diameter of our tree decomposition. We first obtain a binary balanced decomposition using Lemma 3.7, then reindex the vertices in such a way that we can check if a vertex is in some bag of a given subtree of tree decomposition in constant time. Using this, we can in time $k^{O(1)}$ determine whether a bag is a balanced separator of $S$, and if not move to the subtree containing the most vertices of $S$. This procedure will consider at most $d$ bags, hence the total running time of $k^{O(1)}d$. Since the decomposition has depth $O(\log n)$ it also has diameter $d=O(\log n)$. Hence the construction of Wood can be executed in time $k^{O(1)}n\log n$.
@@
To improve on the function of $n$ in the running time of our procedure to compute tree-partitions of width $O(\Delta w)$, we can use some of the techniques introduced in [BDD+16]. If we use separators that are balanced with respect to the subgraph that has to be decomposed, we obtain a balanced decomposition as observed by Reed [Ree92] which gave an approximation algorithm for tree decompositions with running time $f(k)n\log n$. If, in addition, we stop processing once we reach components of size $O(\log n)$, we have computed at most $f(k)n/\log n$ separators which can each be found in time $f(k)\log n$ using the data structure introduced in [BDD+16]. This means that getting to components of size $O(\log n)$ takes time $f(k)n$. On each of the obtained components, we can either apply our previous $f(k)n\log n$ construction, which leads to an $f(k)n\log\log n$ algorithm, or apply our construction recursively to obtain an $f(k)n\log^{(\alpha)}n$ algorithm where $\log^{(\alpha)}$ is the $alpha$-fold composition of $\log$.
@@
A balanced tree-partition of width $O(\Delta w)$ can be computed in time $2^{O(k)}n\log^{*}n$ for a graph of maximum degree $\Delta=k^{O(1)}$ when given a tree decomposition of width $w=O(k)$.
Proof 3.10.
@@
We first observe that to make the decomposition balanced, we only need to add a balanced separator once per bag of the tree-partition which still gives a width of $O(\Delta w)$. For each bag of the constructed tree-partition, we compute a balanced separator once and we compute $O(\Delta)$ $S$-balanced separators. These can be computed in time $2^{O(k)}\log n$ by the data structure after its initialization in time $2^{O(k)}n$. One might worry that because we look at sets $S$ of size polynomial in $k$ and not linear unlike [BDD+16], the data structure does not work or has a running time $2^{O(|S|)}$ instead of $2^{O(k)}$. However, the exponential part in the analysis is an exponential in the width of the tree decomposition. The size of $S$ does have an impact on the running time, but only appears in a polynomial factor. Since we bound the size of $S$ by a polynomial in $k$, the polynomial in $|S|$ still gives a polynomial in $k$ in our setting, and it is still dominated by the exponential in $k$.
@@
We prove the existence of an algorithm of running time $O(2^{O(\alpha)}n\log^{(\alpha)}n)$ for every $\alpha\geq 1$ by induction on $\alpha$.

### v6

@@ Abstract
parameterized algorithms, tree-partitions, tree-partition-width, tree-cut width, domino treewidth, treewidth, approximation algorithms, parameterized complexity
@@ 1 Introduction
Graph decompositions have been a very useful tool to draw the line between tractability and intractability of computational problems. There are many meta-theorems showing that a collection of problems can be solved efficiently if a decomposition of some form is given (e.g. for treewidth [15], for clique-width [16], for twin-width [12], for mim-width [3]). By finding efficient algorithms to compute a decomposition if it exists, we deduce the existence of efficient algorithms even if the decomposition is not given. In particular, this proves useful when designing win-win arguments: for some problems, the existence of a solution and the existence of a decomposition are not independent, so that we can either use the decomposition for an efficient computation of the solution, or conclude that a solution must (or cannot) exist when there is no decomposition of small enough width.
The most successful notion of graph decomposition to date is certainly tree decompositions, and its corresponding parameter treewidth. Any problem expressible in MSO2111Formulae with quantification over sets of edges or vertices, quantification over vertices and edges, and with the incidence predicate. can be solved in linear time in graphs of bounded treewidth due to a meta-theorem of Courcelle [15] and the algorithm of Bodlaender for computing an optimal tree decomposition [4]. Treewidth is a central tool in the study of minor-closed graph classes. A minor-closed graph class has bounded treewidth if and only if it contains no large grid minor.
In this paper, we focus on the parameter tree-partition-width (also called strong treewidth) which was independently introduced by Seese [31] and Halin [26]. It is known to have simple relations to treewidth [17, 33]: $\operatorname{\textup{{tw}}}=O(\operatorname{\textup{{tpw}}})$, and $\operatorname{\textup{{tpw}}}=O(\Delta\operatorname{\textup{{tw}}})$, where $\operatorname{\textup{{tw}}},\operatorname{\textup{{tpw}}},\Delta$ denote the treewidth, the tree-partition-width, and the maximum degree respectively. Applications of tree-partition-width include graph drawing and graph coloring [13, 25, 19, 20, 34, 2, 1]. Recently, Bodlaender, Cornelissen and Van der Wegen [6] showed for a number of problems (in particular, problems related to network flow) that these are intractable (XNLP-complete) when the pathwidth is used as parameter, but become fixed parameter tractable when parameterized by the width of a given tree-partition. This raises the question of the complexity of finding tree-partitions. We show that computing tree-partitions of approximate width is tractable.
@@
Thus, this removes the requirement from the results from [6] that a tree-partition of small width is part of the input. Our technique is modular and allows us to also give alternatives running in FPT time or polynomial time with an improved approximation factor (see Theorem 3.11). Although not formulated as an algorithm, a construction of Ding and Oporowski [18] implies an FPT algorithm to compute tree-partitions of width $f(k)$ for graphs of tree-partition-width $k$, for some fixed computable function $f$. We adapt their construction and give some new arguments designed for our purposes. This significantly improves on the upper bounds to the width, and the running time.
The results from [6] are stated in terms of the notions of stable gonality, stable tree-partition-width and a new parameter called weighted tree-partition-width [^2] In earlier versions of [6], the parameter was called treebreadth, but to avoid confusion, the term weighted tree-partition-width is now used.. The notion of stability comes from the algebraic geometry origins of the notion of gonality; in graph-theoretic terms, this implies that we look at the minimum over all possible subdivisions of edges. It turns out that tree-partition-width, stable tree-partition-width, and weighted tree-partition-width (with edge weights one) are bounded by polynomial functions of each other; see Section 5 and Corollary 5.5. In Section 6, we obtain some results on the complexity of computing and approximating the weighted tree-partition-width, as corollaries of earlier results.
Related to tree-partition-width is the notion of domino treewidth, first studied by Bodlaender and Engelfriet [8]. A domino tree decomposition is a tree decomposition where each vertex is in at most two bags. Where graphs of small tree-partition-width can have large degree, a graph of domino treewidth $k$ has maximum degree at most $2k$. Bodlaender and Engelfriet show that Domino Treewidth is hard for each class $W[t]$, $t\in\mathbb{N}$; we improve this result and show XALP-completeness.
@@
We also consider the parameter tree-cut width introduced by Wollan in [32]. As the tractability results of Bodlaender et al. [6] use techniques similar to a previous work on algorithmic applications of tree-cut width [24], one may wonder whether there is a relationship between tree-cut width and tree-partition-width.
@@
  • We obtain a parameter that is polynomially tied to tree-partition-width and is topological minor monotone (see Theorem 5.4). We use this to show that tree-partition-width is relatively stable with respect to subdivisions: if we define $\underline{\operatorname{\textup{{tpw}}}}(G)$ (resp. $\overline{\operatorname{\textup{{tpw}}}}(G)$) as the minimum (resp. maximum) tree-partition-width over subdivisions of $G$, then $,~{}\operatorname{\textup{{tpw}}}$ tpw and $\overline{\operatorname{\textup{{tpw}}}}$ are polynomially tied. The parameter $\underline{\operatorname{\textup{{tpw}}}}(G)$ corresponds to ‘stable tree-partition-width’.
@@
  • On the other hand, a bound on tree-partition-width does not imply a bound on tree-cut width (see Observation 5.1).
@@ 2 Preliminaries
We also consider a variant of the notion for weighted graphs. The notion was introduced in [6]. We use a slight generalization where we allow also weights of vertices, to facilitate some of our proofs.
@@ 3.1 Description of the algorithm
As mentioned above, we do not directly specify the function $w=w(k)$, since different algorithms for step 1 give different solution qualities (bounds for $w(k)$) and running times. Since $\operatorname{\textup{{tw}}}+1\leq 2\cdot\operatorname{\textup{{tpw}}}$, if $\operatorname{\textup{{tw}}}(G)>2k-1$ it follows that $\operatorname{\textup{{tpw}}}(G)>k$. Suppose that we obtained a decomposition of width $w$ (which could be larger than $\operatorname{\textup{{tw}}}(G)$). If $\operatorname{\textup{{tpw}}}(G)\leq k$, then $\operatorname{\textup{{tw}}}(G)\leq 2k-1$, and so there are at most $(2k-1)n$ edges in $G$. If $G$ has more than $(2k-1)n$ edges, we directly reject.
@@
We later describe several ways of computing the edges of $G^{b}$. We will show in Claim 3.1 that vertices in the same connected component of $G^{b}$ must be in the same bag for any tree-partition of width at most $k$. For this reason, we conclude that $\operatorname{\textup{{tpw}}}(G)>k$ if a component of $G^{b}$ has more than $k$ vertices.
@@
Given the components of $G^{b}$, we can compute $H$, and its 2-connected components in time $O(k^{O(1)}n)$. Using Claim 3.2, we obtain a tree decomposition of $H$ by replacing vertices of $G$, in the tree decomposition of $G$, by their component in $G^{b}$.
By Claim 3.3, the maximum degree $\Delta_{H}$ within the 2-connected components of $H$ is at most $Cbk^{2}$ for some constant $C$ when $\operatorname{\textup{{tpw}}}(G)\leq k$.
@@
By rooting the decomposition of $H$ in 2-connected components, we can define a parent cut-vertex for each 2-connected component except the root. We separately compute tree partitions for each 2-connected component of $H$ with the constraint that their parent cut-vertex should be the single vertex of its bag. A construction of Wood [33] enables us to compute a tree-partition of width $O(\Delta w)$ for any graph of maximum degree $\Delta$ and treewidth $w$; this can be adjusted to allow for this isolation constraint without increasing the upper bound on the width. We give the details of this in Corollary 3.5. After doing this, the partitions of each component can be combined without increasing the width. Indeed, although cut-vertices are shared, only one 2-connected component will consider putting other vertices in its bag. We obtain a tree-partition of $H$ of width $O(wbk^{2})$.
@@ 3.2 Correctness
Assume that $s$ and $t$ are in adjacent bags but not in the same bag for some tree-partition of width at most $k$. We denote their respective bags by $B_{s}$ and $B_{t}$. Then, $(B_{s}\cup B_{t})\setminus\{s,t\}$ is an $s$-$t$ separator of $G-st$. Consequently, $\mu(s,t)\leq 2k-2$. This proves the second point. ∎
Claim 3.2.
@@
Proof.
Every component of $G^{b}$ appears in at least one $Y_{i}$, because it contains a vertex which must appear in at least one $X_{i}$. Furthermore, for each edge $UV$ of $H$, there must be vertices $u\in U,v\in V$ such that $uv$ is an edge of $G$. Hence, there is a bag $X_{i}$ containing $u$ and $v$ so $Y_{i}$ contains $U$ and $V$. Finally, suppose that there is a bag $Y_{i}$ not containing a component $C$ of $G^{b}$, and several components of $T-i$ have bags containing $C$. There must be an edge of $G^{b}$ connecting vertices $u$ and $v$ of $C$ such that $u$ is in bags of $X$ and $v$ is in bags of $X^{\prime}$, where $X$ and $X^{\prime}$ are in different components of $T-i$. By definition of $G^{b}$, the minimal size of a separator of $u$ and $v$ in $G$ is at least $b\geq w+1$. However, since the tree decomposition $(T,(X_{i}))$ has width $w$ and the bags containing $u$ are disjoint from the bags containing $v$ (in particular $X_{i}$ separates them), there is a separator of $u$ and $v$ of size at most $w$, a contradiction. This concludes the proof that $(T,(Y_{i}))$ is a tree decomposition of $H$. ∎
Claim 3.3.
@@
Proof.
Consider $u$ a vertex achieving maximum degree in $B$. By definition of $B$, $B-u$ is connected. We denote by $N$ the neighborhood of $u$ in $B$. Let $T$ be a spanning tree of $B-u$. We iteratively remove leaves that are not in $N$. This produces the reduced tree $T^{\prime}$. The maximum degree in this tree is $b-1$ as the set of edges incident to a given vertex can be extended to disjoint paths leading to vertices in $N$, hence leading to $u$. Since $H$ is a $b$-reduction, the number of disjoint paths between two vertices must be less than $b$.
Clearly the neighbors of $u$ must be either in the same bag as $u$ or in a neighboring bag. Since the bag of $u$ will be a separator of vertices that are in distinct neighboring bags, in particular, it splits the graph into several components each containing at most $k$ neighbors of $u$.
There must exist a subset of vertices of $T^{\prime}$ of size at most $k-1$ whose removal splits $T^{\prime}$ in components containing vertices of $N$ of total weight at most $k$. Since the degree of a vertex of $T^{\prime}$ is at most $b-1$, removing one of its vertices adds at most $b-2$ new components. Hence, after removing $k-1$ vertices, there are at most $1+(k-1)(b-2)$ components. We conclude that $|N|\leq k(1+(k-1)(b-2))$. Since $u$ had maximum degree in $B$, we conclude. ∎
In [33], Wood shows the following lemma.
@@
We deduce this slightly stronger version of [33, Theorem 1]
@@
Proof.
We wish to apply Lemma 3.4 to $S\supseteq N(v)$. Let $\gamma=1+\sqrt{2}$. We have $|N(v)|\leq\Delta$, so in particular, $|N(v)|\leq 3(\gamma+1)(w+1)\Delta$. In case $|N(v)|<(\gamma+1)(w+1)$, we can add arbitrary vertices to $N(v)$ to form $S$ satisfying $|S|\geq(\gamma+1)(w+1)$. Otherwise, we simply set $S=N(v)$. We then apply the lemma to $S$ in $G-v$. There is a single bag that contains $N(v)$, and so we may add the bag $\{v\}$ adjacent to this in order to deduce a tree partition of $G$ of width $O(\Delta w)$ in which $v$ is the only vertex of its bag. ∎
@@ 3.3 Time/quality trade-offs
  • An algorithm of Korhonen [28] computes a tree decomposition of width at most $12k+$ 1 or reports that $\operatorname{\textup{{tw}}}(G)>k$ in time $2^{O(k)}n$.
  • An algorithm of Fomin et al. [22] computes a tree decomposition of width $O(k^{2})$ or reports that $\operatorname{\textup{{tw}}}(G)>k$ in time $O(k^{7}n\log n)$.
  • An algorithm of Feige et al. [21] computes a tree decomposition of width $O(k\sqrt{\log k})$ or reports that $\operatorname{\textup{{tw}}}(G)>k$ in time $O(n^{O(1)})$.
@@
  • We can use a maximum-flow algorithm (e.g. Ford-Fulkerson [23]) to compute for each pair $\{s,t\}$ of vertices of $G$ whether there are at least $b$ vertex disjoint paths from $s$ to $t$, in time $O(bkn)$. To compute a minimum vertex cut, replace each vertex $v$ by two vertices $v_{\text{in}},v_{\text{out}}$ with an arc from $v_{\text{in}}$ to $v_{\text{out}}$. All arcs going to $v$ should go to $v_{\text{in}}$, and all arcs leaving $v$ should leave $v_{\text{out}}$. All arcs are given capacity 1. We may stop the maximum flow algorithm as soon as a flow of at least $b$ was found. Furthermore, we can reduce the number of pairs $\{s,t\}$ of vertices to check to $O(wn)$, as each pair must be contained in a bag due to $b\geq w+1$. This results in a total time of $O(wbkn^{2})$.
  • We can also use dynamic programming to enumerate all possible ways of connecting pairs of vertices that are in the same bag in time $2^{O(w\log w)}n$, which is sufficient to compute $G^{b}$. A state of the dynamic programming consists of the subset of vertices of the bag that are used by the partial solution, a matching on some of these vertices, up to two vertices that were decided as endpoints of the constructed paths, the number of already constructed paths between the endpoints, and two disjoint subsets of the used vertices that are not endpoints, nor in the matching such that we found a disjoint path from the first or second endpoint to them. The bound of $2^{O(w\log w)}$ follows from the fact that this is a subset of the labeled forests on $w$ vertices. We may assume that our tree decomposition is rooted and binary. We first tabulate answers for each subtree of the decomposition by starting from the leaves, and then tabulate answers for each complement of a subtree by starting from the root and, when branching to some child, combining with the partial solutions of the subtree of the other child. By combining tabulated values for subtrees and their complements, we obtain the sought information.
The $b$-reduction $H$ of $G$ and its 2-connected components can be computed in $O(k^{O(1)}n)$ time (see e.g. [27]), since the size of the graph is $O(k^{O(1)}n)$ here.
We will now make use of the following result due to Bodlaender and Hagerup [10]:
@@
When implementing the construction of Wood for 2-connected components of $H$, the running time is dominated by $O(n)$ queries to find a balanced separator with respect to a set $S$ of size $k^{O(1)}$. After a preprocessing in time $O(kn)$, we can do this in time $k^{O(1)}d$ where $d$ is the diameter of our tree decomposition. We first obtain a binary balanced decomposition using Lemma 3.6, then reindex the vertices in such a way that we can check if a vertex is in some bag of a given subtree of tree decomposition in constant time. Using this, we can in time $k^{O(1)}$ determine whether a bag is a balanced separator of $S$, and if not move to the subtree containing the most vertices of $S$. This procedure will consider at most $d$ bags, hence the total running time of $k^{O(1)}d$. Since the decomposition has depth $O(\log n)$, it also has diameter $d=O(\log n)$. Hence, the construction of Wood can be executed in time $k^{O(1)}n\log n$.
@@
To improve on the function of $n$ in the running time of our procedure to compute tree-partitions of width $O(\Delta w)$, we can use some of the techniques introduced in [7]. If we use separators that are balanced with respect to the subgraph that has to be decomposed, we obtain a balanced decomposition as observed by Reed [29] which gave an approximation algorithm for tree decompositions with running time $f(k)n\log n$. If, in addition, we stop processing once we reach components of size $O(\log n)$, we have computed at most $f(k)n/\log n$ separators which can each be found in time $f(k)\log n$ using the data structure introduced in [7]. This means that getting to components of size $O(\log n)$ takes time $f(k)n$. On each of the obtained components, we can either apply our previous $f(k)n\log n$ construction, which leads to an $f(k)n\log\log n$ algorithm, or apply our construction recursively to obtain an $f(k)n\log^{(\alpha)}n$ algorithm where $\log^{(\alpha)}$ is the $\alpha$-fold composition of $\log$.
@@
A balanced tree-partition of width $O(\Delta w)$ can be computed in time $2^{O(k)}n\log^{(\alpha)}n$, for all integers $\alpha\geq 1$, for a graph of maximum degree $\Delta=k^{O(1)}$ when given a tree decomposition of width $w=O(k)$.
Proof.
@@
We first observe that to make the decomposition balanced, we only need to add a balanced separator once per bag of the tree-partition which still gives a width of $O(\Delta w)$. For each bag of the constructed tree-partition, we compute a balanced separator once and we compute $O(\Delta)$ $S$-balanced separators. These can be computed in time $2^{O(k)}\log n$ by the data structure after its initialization in time $2^{O(k)}n$. One might worry that because we look at sets $S$ of size polynomial in $k$ and not linear unlike [7], the data structure does not work or has a running time $2^{O(|S|)}$ instead of $2^{O(k)}$. However, the exponential part in the analysis is an exponential in the width of the tree decomposition. The size of $S$ does have an impact on the running time, but only appears in a polynomial factor. Since we bound the size of $S$ by a polynomial in $k$, the polynomial in $|S|$ still gives a polynomial in $k$ in our setting, and it is still dominated by the exponential in $k$.
@@
We prove the existence of an algorithm of running time $O(2^{O(k)}n\log^{(\alpha)}n)$ for every $\alpha\geq 1$ by induction on $\alpha$.